OpenStudy (anonymous):

Solve. 6-8 (x+3)+5x=5x-(2x-5)+13

jhonyy9 (jhonyy9):

6-8x-24+5x=5x-2x+5+13 -3x-18=3x+18 -36=6x x=-36/6 x=-6

OpenStudy (anonymous):

x=-6

OpenStudy (anonymous):

\[6-8x-24+5x=5x-2x+5+13\] Group everything in x on the left hand side of the equation and everything without x on the right hand side. \[-8x+2x=18+24-6\] add like items together \[-6x=36\] Divide both sides of the equation by -6 \[x=-6\] Substitute -6 for x to check your answer.

OpenStudy (anonymous):

Simplify. 6-8(x+3) +5x